OpenStudy (roseliberty):

solve for x: 0=-2x^2+2x+24

OpenStudy (bmk614):

-2(x2-x-12)=0 Factor out the -2 -2(x+3)(x-4)=0 Factor the trinomial x+3=0 x-4=0 Set both binomials equal to 0 and solve for x x=-3 x=4 x=-3,4

OpenStudy (roseliberty):

THANKS!

OpenStudy (mathmale):

I noted immediately that -2 is a factor common to all of your terms. Divide your equation by 2, term by term. Write your result here, please.